OpenStudy (leozap1):

Fan & medal... 8. Garrett has deposited $519 in a savings account that earns interest at a rate of 4.2% compounded daily. What will the account balance be in 17 years? A. $540.80 B. $1,044.52 C. $1,059.83 D. $2,179.0

OpenStudy (leozap1):

Please teach me how to solve this...

OpenStudy (anonymous):

So the equation for compound interest is A = P(1 + r/n)^nt Each variable represents: A = Accrued Amount (principal + interest) P = Principal Amount r = Annual Nominal Interest Rate as a decimal t = Time Involved in years n = number of compounding periods per unit t; at the END of each period So you plug in the numbers for each variable making the equation A = 519(1+0.042/365)^365*17 Note: The 365 is taken from there being 365 days in a year, since the interest is compounded daily. After putting that into your calculator, you should get C

OpenStudy (leozap1):

p = $519 r = 0.42 t = 17 n = ?

OpenStudy (leozap1):

How did you find n?

OpenStudy (anonymous):

n=365 because that is how many times the interest is compounded in a year

OpenStudy (leozap1):

Oh, okay thank you. I also have this question Rebecca deposited $5,338.18 into a savings account with an interest rate of 5.9% compounded twice a year. About how long will it take for the account to be worth $10,000? Would that mean the equation is "A = 5338.11(1 + .059/2)^2*t"?

OpenStudy (anonymous):

In this problem you know what A is, but not t. So your equation would look like 10,000 = 5,338.18(1+0.059/2)^2*t

OpenStudy (leozap1):

What would be the answer to that question? My answer list is A. 15 years, 11 months B. 32 years, 5 months C. 10 years, 10 months D. 11 years, 0 months The equation gave me "144.99610852".

OpenStudy (leozap1):

If I rounded that to 145 and that represents months that's 12 years and 1 month, but that isn't an answer choice. (Sorry for all the questions.)

OpenStudy (anonymous):

http://www.calculatorsoup.com/calculators/financial/compound-interest-calculator.php I use this online calculator quite a bit to help me out, but the answer I got is 10.8 years which would be C (: